Bugs Bunny has three types of coins in his bunny-bag: nickels, dimes, and pennies. Thirteen of them are nickels, 3/7 of them are pennies, and 20% of them are dimes. How much money does Bugs have in his bunny-bag?
Answer by edjones(7569)

3/7 + 1/5 + x = 1
15+7+35x=35 multiply each side by 35
22+35x=35
35x=13
x=13/35
13 nickels=13/35 of the coins.
There are 13*.05 + 15*.01 + 7*.10
=.65+.15+.70=$1.50
.
Ed